OpenStudy (anonymous):

Given that sin x = p, find an expression, in terms of p, for sec2 x.

hartnn (hartnn):

its sec^2 x or sec (2x) ?

hartnn (hartnn):

if it is sec^2 x then first find cos x from sin x then sec x is just 1/cos x

OpenStudy (anonymous):

can you show me how it works ;; A ;; i just can't get rid of the sinx

OpenStudy (cwrw238):

sin x = p sin^2 x = p^2 cos^2 x = 1 - sin^2x cos^2x = 1 - p^2 sec^2x = 1 / cos^2x = = 1 / ( 1 - p^2)

OpenStudy (cwrw238):

I've assumed you want sec^2 x not sec(2x)